ELECTRONIC SPECTROSCOPY OF Be_{2}: EXPERIMENTAL AND THEORETICAL RESULTS

\underline{L. A. KALEDIN},A. L. KALEDIN, M. C. HEAVEN, Department of Chemistry and Cherry L. Emerson Center for Scientific Computation, Emory University, Atlanta, GA 30322; V. E. BONDYBEY, Institut fur Physikalische und Theoretische Chemie der TU Munchen, W-8046 Garching, Germany.

Low-lying electronic states of Be_{2} have been examined using laser excitation techniques. The dimer was formed by pulsed laser ablating Be into a free-jet expansion. Dimer formation was enhanced by liquid nitrogen cooling of the nozzle assembly.

Dispersed fluorescence spectra were recorded following excitation of various vibrational levels of the B ^{1}Σu^{+} state. These spectra revealed bands of the previously unobserved B ^{1}Σu^{+}⟶A′ ^{1}Π_{g}transition. The term energy (T_{0}=13,942\underline{+}20cm^{−1}) and vibrational interval ΔG_{1/2} =717\underline{+}20cm^{−1}) for the A′ state were determined for the first time.

Potential energy curves and electronic transition moments for Be_{2} were calculated using EOM coupled cluster and MRS-DCI levels of theory with a (12s6p3d2f1g)/[5s4p3d2f1g] basis set. The properties of low-lying singlet, triplet, and quintet states were predicted. The MRSDCI results were found to be in excellent agreement with experimental observations^{a}

aV. E. Bondybey, Chem. Phys. Lett. 109, 436 (1984)